OpenStudy (anonymous):

MEDAL AND FAN AWARDED! Determine which system below will produce infinitely many solutions. -6x + 3y = 18 4x - 3y = 6 2x + 4y = 24 6x + 12y = 36 3x - y = 14 -9x + 3y = -42 5x + 2y = 13 -x + 4y = -6

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

Option choices: -6x + 3y = 18 4x - 3y = 6 2x + 4y = 24 6x + 12y = 36 3x - y = 14 -9x + 3y = -42 5x + 2y = 13 -x + 4y = -6